# RETENTION: Resource-Efficient Tree-Based Ensemble Model Acceleration with Content-Addressable Memory

### v1

@@ II-D Motivation
This work is driven by the challenges identified in the observations and previous work;. While nvTCAM holds great promise for accelerating tree-based model inference in resource-constrained environments, the excessive CAM capacity requirement remains a critical challenge that must be addressed before practical implementation. Therefore, this paper proposes RETENTION, an end-to-end framework that significantly reduces CAM capacity requirement, enabling resource-efficient tree-based ensemble model acceleration.
@@ III-A Overview
nvTCAM enables high-speed, parallel inference for tree-based models, making it a promising solution for acceleration. However, substantial memory consumption and redundancy remain key barriers to real-world deployment, limiting its feasibility in resource-constrained environments. In this section, we introduce RETENTION, an end-to-end framework serving as a crucial component to enable practical and resource-efficient acceleration of tree-based model inference with nvTCAM. It addresses the challenges by (1) minimizing model complexity through purity threshold pruning and (2) enhancing memory utilization by introducing an optimized tree mapping scheme with two data placement strategies, namely occurrence-based double reordering (ODR) and similarity-based path clustering (SPC). Fig. 3 presents an overview of RETENTION, detailing its key components and their interactions. The following sections elaborate on the philosophy and implementation.
@@ III-B Purity Threshold Pruning
As the required memory capacity is correlated to both $\textit{paths}$ and $\textit{unique\_conditions}$, pruning models can effectively reduce CAM capacity requirement. Existing pruning techniques for reducing tree-based ensemble model complexity mainly focus on pre-pruning (since post-pruning algorithms are rarely implemented on ensemble models, as explained in Section II-A), and the algorithms fall into two main categories: (1) structural restriction, which constrains the tree’s shape using parameters such as maximum_depth, and (2) split-based pruning, which prevents further splits when subsequent splits are deemed worthless (e.g., minimum impurity decrease). While both methods reduce model complexity, neither takes into account the class distribution of instances within nodes, potentially leading to significant accuracy degradation. For example, in complex datasets, instances from different classes may remain entangled even at the maximum depth, resulting in serious accuracy loss when forced truncation occurs. Similarly, if an early split does not effectively separate instances, later splits might still provide meaningful differentiation, yet split-based pruning prematurely terminates such opportunities. Although these pruning techniques have minimal impact on boosting-based models, where trees iteratively correct previous errors, bagging-based models can suffer severe accuracy degradation due to their independently trained trees, which lack an error correction mechanism.
@@
To avoid potential catastrophic accuracy degradation, this work introduces purity threshold pruning, a novel pruning algorithm designed for bagging-based models. The workflow is depicted in the left part of Fig. 3. Unlike conventional training, which only takes a dataset as input, our approach incorporates a user-specified tolerance for OOB accuracy loss. During training, each node records its majority class and the corresponding purity (i.e., proportion). Once the ensemble is constructed, rather than concluding with OOB estimation, we iteratively determine the minimum purity threshold that maintains accuracy within the specified tolerance. Nodes exceeding this threshold are converted into leaf nodes and assigned their respective majority class, as illustrated in the right part of Fig. 3.
The core principle of purity threshold pruning is to reduce model complexity while preserving accuracy at the ensemble level. Since individual trees are trained on different subsets with distinct splits, a minority class overlooked in one tree can still be classified correctly in others. For highly imbalanced datasets, applying class weighting helps prevent the model from favoring majority classes. Additionally, by explicitly considering OOB accuracy during pruning, the method ensures that performance degradation remains controlled. Unlike standard pruning techniques that may create leaf nodes with evenly distributed classes, potentially leading to misclassifications, purity threshold pruning ensures that each leaf node maintains sufficient class purity to guarantee the desired level of accuracy. Moreover, when early splits already satisfy the target accuracy, further splits are omitted, significantly reducing both $\textit{paths}$ and $\textit{unique\_conditions}$ compared to conventional pruning methods. Different from most post-pruning algorithms that operate on individual decision trees in isolation, purity threshold pruning is specifically designed to account for the collective behavior of ensemble models during inference. This global perspective enables purity threshold pruning to preserve diversity and effectively reduce model complexity while explicitly constraining accuracy degradation within a user-defined tolerance, a capability not addressed by conventional single-tree pruning methods.
Furthermore, recent pruning techniques focus on removing entire trees from models if they exhibit substantial redundancy or structural resemblance [44, 45]. Since purity threshold pruning operates at the node level and assumes trees are diversely grown, it is fully compatible with such whole-tree pruning methods. This compatibility allows further optimization in space efficiency without sacrificing accuracy, making purity threshold pruning a highly adaptable approach for resource-efficient tree-based model. However, the algorithm is not compatible with boosting-based models, as performing pruning after model construction destroys the error correction mechanism, leading to catastrophic accuracy degradation. Nonetheless, the pre-pruning algorithms can already sufficiently reduce boosting-based model complexity without compromising accuracy.
@@
Width per Unit $\left\lceil\frac{\#\textit{unique\_conditions}}{S}\right\rceil$ TCAMs $\left\lceil\frac{\#\textit{unique\_conditions}}{S}\right\rceil$ TCAMs
Height per Unit $\left\lceil\frac{\#\textit{paths}}{S}\right\rceil$ TCAMs $\left\lceil\frac{\#\textit{paths}}{S}\right\rceil$ TCAMs
@@
Key Strength lower input pre-processing cost higher space efficiency
@@ III-C Input pre-processing during Inference
To accelerate tree-based model inference with CAM, raw input data must be transformed into the specific query format for individual search operations. This transformation involves two key steps: (1) feature encoding and (2) query packing. For feature encoding, the thresholds of each feature must be sorted before inference. When raw input data for a feature is received, binary search can be employed to accelerate condition checks. This step can also be performed on the sensor side, as modern smart sensors are capable of handling such lightweight tasks [46]. Additionally, since each feature is typically associated with a dedicated sensor, pre-processing at the sensor level is feasible and efficient. Once features are encoded into binary sequences, the query packing step organizes condition check results into the query format, ensuring alignment with the order of TCAM columns. These queries are then transmitted to the corresponding TCAM via a network-on-chip (NoC) [47], following the same implementation as in [24]. The input pre-processing for ACAM is almost the same, as quantization can be interpreted as a sequence of condition checks. However, since ACAM performs search operations using analog values, additional digital-to-analog conversions are required after receiving the query.
While tree traversal is essentially a sequence of condition checks, encoding input features into sequences and performing in-memory search may initially seem redundant. However, due to the shared conditions across the ensemble, the actual number of unique conditions is significantly lower than expected. Moreover, binary search can further reduce the number of condition checks. For instance, in a Random Forest model with 100 trees and an average path length of 17.38, traversing the forest requires about 1738 condition checks. However, experimental results reveal that as there are 5446 unique conditions within the model, applying binary search can reduce the number of condition checks to approximately 120 ($14~\textit{features}\times\log_{2}\frac{5446}{14}$), which is a significant improvement. Additionally, feature encoding is highly efficient compared to conventional CPU-based tree traversal, because all data accesses are deterministic and easily cached. Given that the computational overhead for feature encoding is constant when using TCAM as an accelerator, and can also be done on the sensor side, this work focuses on the computational overhead of query packing, which is primarily determined by the selected data placement strategy.
@@ III-D1 Naive Tree Mapping Approaches
To utilize TCAM for acceleration, each root-to-leaf path within the model needs to be mapped to a TCAM row, with every column of the TCAM corresponding to a unique condition within the paths being searched. There are two naive approaches for mapping: (1) naive unified mapping and (2) naive independent mapping, with their differences summarized in Table II, assuming a TCAM size of $\textit{S}\times\textit{S}$. In naive unified mapping, paths from the entire ensemble are mapped together, so that each row entry corresponds to the same set of conditions. This structure allows for efficient input pre-processing since TCAMs in the same column receive identical input sequences. However, most nodes in the ensemble are not encountered by any single path, resulting in a vast number of cells storing the X state. In fact, mapping 100 trees together can lead to over 96% redundancy, as shown in Table I. Conversely, naive independent mapping maps each tree separately, ensuring that conditions unique to a tree do not introduce unnecessary columns. This approach significantly reduces storage redundancy, as only the conditions relevant to a specific tree are stored, making it far more space-efficient than naive unified mapping. However, this independence disrupts input pre-processing because each TCAM now requires a distinct input sequence, increasing pre-processing overhead for query packing.
Although naive unified mapping generally demands more TCAMs than naive independent mapping, the energy-saving precharge mechanism proposed in DT2CAM [22] can help mitigate energy waste by deactivating unmatched rows early. Moreover, input pre-processing dominates energy consumption in inference acceleration, as CPU computations are significantly more power-intensive than in-memory search operations. Given that many nodes are shared among multiple trees, naive unified mapping reduces input pre-processing overhead, making it more energy-efficient than naive independent mapping. Therefore, based on the above observations, we categorize the optimization criteria of data placement strategies into two types: energy-efficient mapping and space-efficient mapping. The key difference between these two types is whether to minimize memory usage at the cost of increased computational overhead.
Figure 4: Visualization of the proposed data placement strategies.
@@ III-D2 Existing Data Placement Strategy Analysis
DT2CAM primarily focused on mapping a single decision tree to TCAM without addressing redundancy or ensemble models. When directly applied to ensembles, DT2CAM acts as either naive unified mapping or naive independent mapping, depending on the chosen mapping unit. Pedretti et al. [23] introduced a data placement strategy, referred to as the feature reordering with row elimination (FR) algorithm, for ACAM. This method selects the whole ensemble as a mapping unit, and proposes (1) reordering features based on frequency to cluster redundant cells, and (2) eliminating entire path segments when they become redundant. However, due to row elimination, paths were no longer stored in fixed rows, making the result retrieval process substantially more complex and CPU-dependent. As a result, the additional computational overhead is proportional to $\textit{ACAMs}$, similar to that of naive independent mapping. Thus, we classify FR as a space-efficient mapping approach, though its space efficiency may still fall short of naive independent mapping. A follow-up study by Pedretti et al. [24] proposed fitting each tree into a separate ACAM without mitigating redundancy, resembling naive independent mapping. Consequently, we also categorize this approach under space-efficient mapping.
@@ III-D3 Occurrence-Based Double Reordering (ODR)
4:paths $\leftarrow\{\}$
5:sorted_conditions $\leftarrow$ conditions.sort() $\triangleright$ descending freq.
6:for c in sorted_conditions.reverse() do
@@
11:return sorted_conditions, paths
To alleviate memory redundancy in an energy-efficient manner, each path should be stored in a fixed row to avoid additional computational overhead during result retrieval. Moreover, the input sequences for TCAMs in the same column must remain consistent to minimize input pre-processing overhead for query packing. Given these constraints, we propose ODR to enhance memory utilization for resource-constrained environments that prioritize energy efficiency. As outlined in Algorithm 1, ODR first sorts conditions by occurrence frequency in descending order, similar to FR. Instead of eliminating partial sequences when an entire TCAM row is irrelevant to a given path, which introduces sequential computational overhead as seen in FR, ODR optimizes path ordering based on condition usage. Paths that contain rare conditions are placed at the top, concentrating most of the X-state cells in the bottom-right TCAMs. TCAMs that are entirely filled with X-state cells can then be removed, reducing redundancy without distributing a single path across multiple rows, thereby avoiding additional computational overhead.
Despite its effectiveness, ODR still leaves considerable redundancy, as shown in Fig. 4. For example, in the top-right TCAM after applying ODR, only a single meaningful cell remains, yet the entire TCAM is required for correct functionality. To further improve memory efficiency, the next section introduces a more aggressive data placement strategy that reduces CAM capacity requirement at the cost of higher energy consumption. However, its computational overhead is comparable to that of naive unified mapping, and is, in fact, even lower in practice.
@@ III-D4 Similarity-Based Path Clustering (SPC)
Since memory redundancy arises from unencountered conditions, clustering paths with high similarity (i.e., those that share many conditions) can substantially improve space efficiency. Although naive independent mapping can be viewed as a basic form of clustering, paths within a tree exhibit limited similarity, and some TCAMs can still remain nearly redundant especially in cases where $\textit{paths}=\textit{S}+1$ and the TCAM size is $\textit{S}\times\textit{S}$, leading to inefficient capacity utilization. Finding the optimal clusters to minimize TCAM consumption is an NP-hard problem. In this section, we propose SPC, a heuristic algorithm that greedily maximizes similarity among paths within TCAM, and thus minimizes CAM capacity requirement. As presented in Algorithm 2, SPC starts with a pool containing all root-to-leaf paths in the model. While pool is not empty, SPC evaluates the similarity between current_cluster and each of the paths within pool, and designates the path that can fit in current_cluster based on (1) the highest number of shared conditions, and (2) the smallest resulting set of unique conditions after being integrated into current_cluster, as candidate. Since the objective is to accommodate all paths using the smallest number of clusters, and the capacity of each cluster depends on both $\textit{paths}$ and $\textit{unique\_conditions}$ because we limit a cluster to a single TCAM, the key intuition behind SPC is to minimize $\textit{unique\_conditions}$ at every step. This approach allows each TCAM to hold as many paths as possible, resulting in an effective heuristic. Once current_cluster reaches capacity or none of the paths is suitable, current_cluster is added to clusters representing the completed ones, and a new empty cluster is initiated. Otherwise, SPC greedily adds the candidate to current_cluster and removes it from pool, and this process is repeated until all paths have been assigned to clusters.
SPC efficiently exploits similarities across paths, diminishing unnecessary X states. By filling TCAMs as fully as possible, it maximizes utilization and minimizes wasted capacity. Since each cluster is constrained to contain at most S unique conditions and S paths, there is no dependency between the match results of different TCAMs, eliminating any computational overhead associated with intersecting partial match results. Therefore, the only computational overhead lies in the query packing for every TCAM, which is the same as naive independent mapping. However, since SPC requires far fewer TCAMs than naive independent mapping, it emerges as a significantly more resource-efficient approach in terms of both energy and memory consumption.
@@
4:clusters $\leftarrow\{\}$, current_cluster $\leftarrow\{\}$
5:current_num $\leftarrow 0$
6:while pool is not empty do
7:  similarity $\leftarrow$ calc_similarity(pool, current_cluster)
8:  candidate $\leftarrow$ find_best_candidate(similarity, pool)
9:  if candidate is NULL or current_num == S then
10:   clusters.append(current_cluster)
11:   current_cluster $\leftarrow\{\}$
12:   current_num $\leftarrow 0$
13:  else
14:   current_cluster.append(candidate)
15:   current_num $\leftarrow$ current_num + 1
16:   pool.remove(candidate)
17:if current_cluster is not empty then
18:  clusters.append(current_cluster)
19:return clusters

### v2

@@ II-D Motivation
This work is driven by the challenges identified from the previous works mentioned above. While nvTCAM holds great promise for accelerating tree-based model inference in resource-constrained environments, the excessive CAM capacity requirement remains a critical challenge that must be addressed before practical implementation. Therefore, this paper proposes RETENTION, an end-to-end framework that significantly reduces CAM capacity requirement, enabling resource-efficient tree-based ensemble model acceleration.
@@ III-A Overview
nvTCAM enables high-speed, energy-efficient inference for tree-based models, making it a promising solution for acceleration. However, substantial memory consumption and redundancy remain key barriers to real-world deployment, limiting its feasibility in resource-constrained environments. In this section, we introduce RETENTION, an end-to-end framework serving as a crucial component to enable practical and resource-efficient acceleration of tree-based model inference with nvTCAM. It addresses the challenges by (1) minimizing model complexity through purity threshold pruning and (2) enhancing memory utilization by introducing an optimized tree mapping scheme with two data placement strategies, namely occurrence-based double reordering (ODR) and similarity-based path clustering (SPC). Fig. 3 presents an overview of RETENTION, detailing its key components and their interactions. The following sections elaborate on the philosophy and implementation of each component, concluding with a discussion on the generalizability of RETENTION to ACAM.
@@ III-B Purity Threshold Pruning
As the required CAM capacity is correlated to both $\textit{paths}$ and $\textit{unique\_conditions}$, pruning models can achieve effective mitigation by reducing both at the same time. Existing pruning techniques for reducing tree-based ensemble model complexity mainly focus on pre-pruning (as explained in Section II-A), and the algorithms fall into two main categories: (1) structural restriction, which limits the tree’s shape with hard constraints (e.g., maximum depth), and (2) split-based pruning, which stops splitting when subsequent splits are deemed worthless (e.g., minimum impurity decrease). While both methods reduce model complexity, neither takes into account the class distribution of instances within nodes, potentially leading to significant accuracy degradation. For example, in complex datasets, instances from different classes may remain entangled even at the maximum depth, degrading accuracy substantially when forced truncation occurs. Similarly, if an early split does not effectively separate instances, later splits might still provide meaningful differentiation, yet split-based pruning prematurely terminates such opportunities. While these techniques have minimal impact on boosting-based models, where trees iteratively correct previous errors, bagging-based models can suffer severe accuracy loss due to their independently trained trees, which lack an error correction mechanism.
@@
To avoid potential catastrophic accuracy degradation, this work introduces purity threshold pruning, a novel pruning algorithm designed for bagging-based models. The workflow is depicted in the left part of Fig. 3. Unlike conventional training, which only takes a dataset as input, our approach incorporates a user-specified tolerance for OOB accuracy loss. During training, each node records its majority class and the corresponding purity (i.e., proportion). Once the ensemble is constructed, rather than concluding with OOB estimation, we iteratively determine the minimum purity threshold that maintains accuracy within the specified tolerance. Nodes exceeding this threshold are converted into leaf nodes representing their respective majority class, as shown in the right part of Fig. 3.
@@
Key Strength lower input pre-processing cost lower capacity requirement
@@
The core principle of purity threshold pruning is to reduce model complexity while preserving accuracy at the ensemble level. Since individual trees are trained on different subsets of training data with distinct splits, a minority class overlooked in one tree can still be classified correctly in others. For highly imbalanced datasets, applying class weighting helps prevent the model from favoring majority classes. Additionally, by explicitly considering OOB accuracy during pruning, the method ensures that performance degradation remains controlled. Unlike standard pruning techniques that may create leaf nodes with evenly distributed classes, potentially leading to misclassifications, purity threshold pruning ensures that each leaf node maintains sufficient class purity to guarantee the desired level of accuracy. Moreover, when early splits already satisfy the target accuracy, further splits are omitted, significantly reducing both $\textit{paths}$ and $\textit{unique\_conditions}$ compared to conventional pruning methods. Different from most post-pruning algorithms that operate on individual decision trees in isolation, purity threshold pruning is specifically designed to account for the collective behavior of ensemble models during inference. This global perspective enables purity threshold pruning to preserve diversity and effectively reduce model complexity while explicitly constraining accuracy degradation within a user-defined tolerance, a capability not possessed by conventional single-tree pruning methods.
Recent innovations in pruning focus on removing entire trees from models if they show substantial redundancy or structural resemblance [45, 47]. Since purity threshold pruning operates at the node level and assumes trees are diversely grown, it is fully compatible with such whole-tree pruning methods, exhibiting high adaptability for resource-constrained environments. However, it is not compatible with boosting-based models, as pruning such models after construction ruins their error correction mechanism, leading to catastrophic accuracy degradation. Still, the sequential training of boosting-based models allow existing pre-pruning algorithms to sufficiently reduce complexity without compromising accuracy.
@@ III-C Input Pre-Processing during Inference
To accelerate tree-based model inference with CAM, raw input data must be transformed into the specific query format for individual search operations. This transformation involves two key steps: (1) feature encoding and (2) query packing. For feature encoding, the thresholds of each feature must be sorted in advance for efficient inference. When raw input data for a feature is received, binary search can be employed to accelerate threshold comparisons. This step can be performed on the sensor side in the context of near-sensor inference, as modern smart sensors are capable of handling such lightweight tasks [37]. Additionally, since each feature is typically associated with a dedicated sensor, pre-processing at the sensor level is feasible and efficient. Once features are encoded into binary sequences, the query packing step organizes condition check results into the query format, ensuring alignment with the order of TCAM columns. These queries are then transmitted to the corresponding TCAM via a network-on-chip (NoC) [5], following the same implementation as in [32].
Since tree traversal is essentially a sequence of condition checks, encoding input features into sequences and performing in-memory search may initially seem redundant. However, due to the frequent condition sharing across the ensemble, the actual number of unique conditions is significantly lower than expected. Moreover, binary search can further reduce the number of condition checks for feature encoding. For instance, in a Random Forest model with 100 trees, 14 input features, and an average path length of 17.38, traversing the forest with conventional processors requires about 1738 condition checks. On the other hand, experimental result reveals that there are 5446 unique conditions within the model. Applying binary search can reduce the number of condition checks to approximately 120 ($14~\textit{features}\times\log_{2}\frac{5446}{14}$), which is a significant improvement. Additionally, feature encoding is highly efficient compared to conventional CPU-based tree traversal, as all data accesses are deterministic and easily cached. Given that the computational overhead for feature encoding is constant when using TCAM as an accelerator, this work focuses on the computational overhead of query packing, which is determined by the selected data placement strategy.
Figure 4: Visualization of the proposed data placement strategies.
@@ III-D1 Naive Tree Mapping Approaches
To utilize TCAM for acceleration, each root-to-leaf path within the model needs to be mapped to a TCAM row, with every column of the TCAM corresponding to a unique condition within the paths being searched. There are two naive approaches for mapping: (1) naive unified mapping and (2) naive independent mapping, with their differences summarized in Table II, assuming a TCAM size of $\textit{S}\times\textit{S}$. In naive unified mapping, paths from the entire ensemble are mapped together, so that each row entry corresponds to the same set of conditions. This structure allows for efficient input pre-processing since TCAMs in the same column receive identical input queries. However, most nodes in an ensemble are only encountered by few paths, resulting in a vast number of cells storing the X state. In fact, mapping 100 trees together can lead to over 96% redundancy, as shown in Table I. Conversely, naive independent mapping maps each tree separately, ensuring that conditions unique to one tree do not introduce redundancy to rows associating with paths in other trees. Although this approach significantly alleviates redundancy, each TCAM requires a distinct input query, increasing pre-processing overhead.
While naive unified mapping generally demands more TCAMs than naive independent mapping, the energy-saving precharge mechanism proposed in DT2CAM [35] can help mitigate energy waste by deactivating mismatched rows early. Furthermore, given that input pre-processing dominates energy consumption, as CPU computation is considerably more power-intensive than in-memory search operation (refer to Table IV for detailed statistics), and that naive unified mapping requires fewer iterations of query packing due to query sharing, it is more energy-efficient than naive independent mapping. Therefore, based on the above observations, we categorize the optimization criteria of data placement strategies into two types: energy-efficient mapping and space-efficient mapping. The key difference is whether to minimize capacity requirement at the cost of increased computational overhead.
@@ III-D2 Existing Data Placement Strategy Analysis
DT2CAM primarily focused on mapping a single decision tree to TCAM without addressing redundancy or ensemble models. When directly applied to ensembles, DT2CAM acts as either naive unified mapping or naive independent mapping, depending on the chosen mapping unit. Pedretti et al. [31] introduced a data placement strategy, referred to as the feature reordering with row elimination (FR) algorithm, for ACAM. This method selects the whole ensemble as a mapping unit, and proposes (1) reordering features based on frequency to cluster redundant cells, and (2) eliminating entire path segments in each ACAM when they are fully redundant. However, due to row elimination, paths were no longer stored in fixed rows across different ACAMs, making the merging of partial match results substantially more complex and CPU-dependent. As a result, the additional computational overhead is proportional to $\textit{ACAMs}$, similar to that of naive independent mapping for query packing. Thus, we classify FR as a space-efficient mapping approach, though its capacity requirement may be higher than naive independent mapping. A follow-up study by Pedretti et al. [32] proposed fitting each tree into separate ACAMs, which is identical to naive independent mapping.
@@ III-D3 Occurrence-Based Double Reordering (ODR)
To alleviate memory redundancy in an energy-efficient manner, each path should be stored in a fixed row to avoid additional computational overhead when merging partial match results. Moreover, input queries for TCAMs in the same column must remain consistent to minimize input pre-processing overhead for query packing. Given these constraints, we propose ODR to enhance memory utilization for resource-constrained environments that prioritize energy efficiency. As outlined in Algorithm 1, ODR first sorts conditions by occurrence frequency in descending order, similar to FR. Instead of eliminating partial sequences when a TCAM row is entirely filled with X states, which introduces sequential computational overhead as seen in FR, ODR optimizes path ordering based on condition usage. Paths that contain rare conditions are placed at the top, concentrating most of the cells storing the X state in the bottom-right TCAMs. TCAMs that are utterly filled with X states can then be removed, reducing redundancy without additional computational overhead.
Despite its effectiveness, ODR still leaves considerable redundancy, as shown in Fig. 4. For example, there is only one meanful cell in the top-right TCAM after applying ODR, yet the entire TCAM is required for correct functionality. To further reduce capacity requirement, the next section introduces a more aggressive data placement strategy that reduces redundancy at the cost of slightly higher energy consumption. However, its computational overhead is comparable to that of naive unified mapping, and is, in fact, even lower in practice.
@@
4:paths $\leftarrow[\ ]$
5:sort(conditions, descending)
6:for c in conditions.reverse() do
@@
11:return conditions, paths
@@ III-D4 Similarity-Based Path Clustering (SPC)
Since memory redundancy arises from unencountered conditions, clustering paths with high similarity (i.e., those that share many conditions) can substantially reduce capacity requirement. Although naive independent mapping can be viewed as a basic form of clustering, paths within a tree exhibit limited similarity, and some TCAMs can still remain nearly redundant especially in cases where $\textit{paths}=\textit{S}+1$ and the TCAM size is $\textit{S}\times\textit{S}$, leading to inefficient capacity utilization. Finding the optimal clusters to minimize TCAM consumption is an NP-hard problem. In this section, we propose SPC, a heuristic algorithm that greedily maximizes similarity among paths within each TCAM, and thus minimizes capacity requirement. As presented in Algorithm 2, SPC starts with a pool containing all root-to-leaf paths in the model. While pool is not empty, SPC evaluates the similarity between current_cluster and each of the paths within pool, and designates the path that can fit in current_cluster based on (1) the highest number of shared conditions, and (2) the smallest resulting set of unique conditions after being integrated into current_cluster, as candidate. Since the objective is to accommodate all paths using the smallest number of clusters, and the capacity of each cluster depends on both $\textit{paths}$ and $\textit{unique\_conditions}$ because we limit a cluster to a single TCAM, the key intuition behind SPC is to minimize $\textit{unique\_conditions}$ at every step. This approach allows each TCAM to hold as many paths as possible, resulting in an effective heuristic. Once current_cluster reaches capacity or none of the paths is suitable, current_cluster is added to clusters representing the completed ones, and a new empty cluster is initiated. Otherwise, SPC greedily adds the candidate to current_cluster and removes it from pool, and this process is repeated until all paths have been assigned to clusters.
SPC efficiently exploits similarity across paths, diminishing unnecessary X states. By filling TCAMs as fully as possible, it maximizes utilization and minimizes wasted capacity. Since each cluster is constrained to contain at most S unique conditions and S paths, there is no dependency between the match results of different TCAMs, eliminating any computational overhead associated with merging partial match results. Therefore, the only computational overhead lies in the query packing for every TCAM, which is the same as naive independent mapping. However, since SPC requires far fewer TCAMs than naive independent mapping, it emerges as a significantly more resource-efficient approach in terms of both energy and memory consumption.
@@
4:clusters $\leftarrow[\ ]$, curr_cluster $\leftarrow[\ ]$
5:while pool is not empty do
6:  similarity $\leftarrow$ calc_similarity(pool, curr_cluster)
7:  candidate $\leftarrow$ find_best_candidate(similarity, pool)
8:  if candidate is None or curr_cluster.size() == S then
9:   clusters.append(curr_cluster)
10:   curr_cluster $\leftarrow\{\}$
11:  else
12:   curr_cluster.append(candidate)
13:   pool.remove(candidate)
14:if curr_cluster is not empty then
15:  clusters.append(curr_cluster)
16:return clusters
III-E Generalizability of RETENTION to ACAM
III-E1 Input Pre-Processing during Inference
The input pre-processing for ACAM is functionally equivalent to the TCAM flow, as quantization can also be viewed as a sequence of condition checks. However, since ACAM operates on analog values, digital-to-analog conversion is required once the input query is dispatched to the respective ACAM.
III-E2 Tree Mapping Scheme
The fundamental difference between mapping to TCAM and ACAM lies in the meaning of a column. Since each column represents a feature rather than a condition in ACAM, the condition reordering of ODR should be changed to feature reordering, while the similarity of SPC should denote the number of shared features. While RETENTION supports ACAM, the relative reduction in capacity requirement is less pronounced when #features\#features and the model size are both small. This is because ACAM inherently requires fewer columns than TCAM for a given model, leading to diminishing returns for smaller-scale configurations.